OpenStudy (anonymous):

2x+4y-z=7 2x-4y+2z=-6 x+4y+z=0

jhonyy9 (jhonyy9):

assum first with second 2x+4y-z=7 2x-4y+2z=-6 --------------- 4x 0 +z=1 and assum second with 3rd 2x-4y+2z=-6 x+4y+z =0 -------------- 3x+ 0 3z=-6 4x+z=1 multiply by -3 3x+3z=-6 -12x-3z=-3 3x+3z=-6 ------------- assum -9x 0 =-9 -9x=-9 x= -9/(-9) =1 x=1 4(1)+z=1 so 4+z=1 so z=-3 x+4y+z=0 1+4y-3=0 4y=2 y=2/4 = 1/2 x=1 y=1/2 z=-3